What are the values of x that g(x)=0 where g(x)= (x-1)/(x-4) ?
Thank you!

Answer by jim_thompson5910(28595)

g(x)= (x-1)/(x-4)


0 = (x-1)/(x-4)


0(x-4) = x-1


0 = x-1


x - 1 = 0


x = 1


So the solution is x = 1